# Web 3.0 and Quantum Security: Long-Distance Free-Space QSDC for Global Web 3.0 Networksthanks: These authors contributed equally: Yew Kee Wong, *Yifan Zhou, Xinlin Zhou, Yan Shing Liang, and Zi Yan Li

### v3

@@ I Introduction
Communication Distance Long-distance (intercontinental) Moderate distance Moderate distance Short to moderate distance
Security Level High (no key exchange required) High (can transmit with or without key exchange) High (can transmit with or without key exchange) High (key exchange fundamental)
@@
In the evolution of satellite-based QKD, significant strides have been achieved since the launch of the Micius satellite in 2017, a pioneering endeavor that demonstrated the transmission of secure quantum keys over intercontinental distances [3].
Despite these advancements, satellite-based QKD faces persistent challenges, including substantial signal loss, atmospheric interference, and the need for precise alignment and tracking of satellite signals[4, 5, 6].
@@
The DL04 protocol contributes to these advancements by facilitating secure quantum communication without the need for a shared secret key and bypassing the necessity for quantum memory. This protocol ensures the integrity of quantum communication through immediate state preparation and encoding, followed by direct transmission to the receiver. This process is detailed as follows[7]:
State Preparation and Encoding: Similar to other QSDC protocols, the memory-free DL04 requires the sender, Alice, to prepare and encode quantum states. However, instead of storing these states, they are immediately transmitted once prepared.
Quantum State Transmission: Upon preparation, the quantum states are sent directly to the receiver, Bob, who must immediately measure them using predetermined settings agreed upon in advance. This immediate measurement is crucial as it does not allow for the storage and later retrieval of quantum information, thus eliminating the need for quantum memory.
Eavesdropping Detection: The protocol incorporates mechanisms for detecting eavesdropping through quantum state disturbance. Since measurements are performed immediately after transmission, any disturbance caused by an eavesdropper would be detected as anomalies in the measurement results. This approach leverages the fundamental principle of quantum mechanics—any attempt to measure a quantum state will invariably alter it.
Direct Information Transfer: Following successful transmission and measurement, the information is decoded directly from the quantum states. This step confirms the secure transfer of information without the need for intermediate storage, thus negating the requirement for quantum memory.
@@ II Long-Distance Free-Space QSDC Overview
This section introduces long-distance free-space quantum secure direct communication (QSDC). Our investigation delves into how advanced technologies can sur- mount existing constraints, rendering QSDC viable for widespread application both in the open air and across space through satellite communication.
The process initiates by setting up the quantum state, marking the first phase in encoding quantum information for transmission. Following this, the process involves applying LDPC coding to enhance error correction capabilities, crucial for maintaining the integrity of quantum data over long distances. PAT technologies are integrated to ensure precise alignment and stabilization of the quantum signal. An atmospheric correction step is then applied to mitigate effects like scattering and absorption that can degrade the quantum signal as it traverses the atmosphere. The corrected signal undergoes transmission, sending it across free space, potentially covering vast distances including satellite-to-ground communication paths. Upon reaching the destination, the quantum signal is measured in a critical phase where the encoded quantum information is detected and interpreted. Finally, the process concludes with information decoding, where the quantum data is translated back into classical information for use.
@@ II-A Quantum State Preparation and Encoding
Basis Determination ($a_{0}$): - If $a_{0}=0$, the basis is computational, with states $\{|0\rangle,|1\rangle\}$. - If $a_{0}=1$, the basis is superpositional, where $a_{1}$ further dictates the specific state: - $a_{1}=0$: State becomes $\frac{|0\rangle+|1\rangle}{\sqrt{2}}$ $a_{1}=1$: State becomes $\frac{|0\rangle-|1\rangle}{\sqrt{2}}$
@@
- If $a_{2}=0$, the phase states remain as initially encoded. - If $a_{2}=1$, the phase states are swapped: - $|0\rangle+|1\rangle$ becomes $|0\rangle-|1\rangle$ and vice versa.
@@ III-A1 LDPC Code Parameter Optimization
We propose a GNN-enhanced adaptive decoding algorithm for quantum communication, improving LDPC code decoding by adapting to channel changes and enhancing security with decoy states for eavesdropping detection and privacy amplification methods to secure fi- nal keys, significantly increasing transmission reliability and protection. Standard BP Equation[25]:
@@ III-A3 Security Enhancement Techniques
  • After decoding, apply a hash function $H$ to the key to reduce its length and eliminate partial information.
@@ V-A Technical Implementation of LF QSDC
Figure 4: Summarizes the main characteristics of free-space channel and their impacts on optical signal transmission[30]
@@ VIII Acknowledgement
Author Contributions: Under the guidance of Yew Kee Wong, Yifan Zhou, Yan Shing Liang, Xinlin Zhou, and Zi Yan Li conducted the research with Yifan Zhou and Yan Shing Liang leading the methodology development. The original draft was primarily written by Yifan Zhou, with significant contributions from Xinlin Zhou, Zi Yan Li, and Yan Shing Liang. Xinlin Zhou also undertook the task of reviewing and editing the manuscript. Yew Kee Wong supervised the project, provided revisions, and managed project administration.

### v4

# Novel Long Distance Free Space Quantum Secure Direct Communication for Web 3.0 Networksthanks:

@@ I Introduction
Communication Distance Long-distance (intercontinental) Moderate distance Moderate distance Moderate distance
Security Level High (no key exchange required) High (can transmit secure messages without key exchange) High (can transmit secure messages without key exchange) High (key exchange fundamental)
@@
In the evolution of satellite-based QKD, significant strides have been achieved since the launch of the Micius satellite in 2017, a pioneering endeavor that demonstrated the transmission of secure quantum keys over intercontinental distances[3]. Despite these advancements, satellite-based QKD faces persistent challenges, including substantial signal loss, atmospheric interference, and the need for precise alignment and tracking of satellite signals[4, 5, 6].
Challenges with QKD for Global Scale Communication: QKD systems experience significant signal loss and attenuation over long distances, limiting practical distances without quantum repeaters to a few hundred kilometers. Atmospheric disturbances further exacerbate signal degradation in free-space QKD, making it highly susceptible to environmental factors. For instance, substantial signal loss and atmospheric interference hinder the efficiency of QKD in long-distance communications[7, 8].
Extending QKD reach beyond terrestrial limits with satellites introduces additional challenges, including precise satellite alignment and significant signal loss during transmission. Implementing a global QKD network requires substantial infrastructure, including numerous satellites and ground stations, and the integration of quantum repeaters. The high cost and complexity of this infrastructure pose barriers to widespread adoption[9, 10].
Advantages of QSDC for Global Scale Communication In contrast, QSDC protocols, particularly our solution LF QSDC, offer more practical solutions for global communication. QSDC allows direct transmission of secure messages without prior key exchange, reducing communication complexity and eliminating intermediate key management. Additionally, QSDC protocols incorporate eavesdropping detection through quantum state disturbance, ensuring real-time detection of interference[11, 12]. Technologies like Quantum-Aware Low-Density Parity-Check (LDPC) coding, Pointing, Acquisition, and Tracking (PAT) systems, and Atmospheric Quantum Correction Algorithms (AQCA) improve LF QSDC reliability over long distances. These technologies address key challenges in satellite-based free-space communication, ensuring precise alignment and minimizing signal loss due to atmospheric disturbances. LF QSDC’s reliance on advanced error correction and adaptive technologies reduces the need for extensive infrastructure compared to QKD, making it more scalable and cost-effective. By leveraging existing satellite and communication infrastructure with minimal modifications, LF QSDC can be integrated seamlessly into current systems[13, 14].
Thus, while QKD has laid the groundwork for secure quantum communication, its limitations in signal loss, infrastructure requirements, and susceptibility to environmental factors make it less suitable for global-scale applications. QSDC protocols, particularly LF QSDC, present a more robust and practical solution for long-distance secure communication, offering enhanced security, scalability, and cost-effectiveness.
@@
The memory-free DL04 protocol contributes to these advancements by facilitating secure quantum communication without the need for a shared secret key and bypassing the necessity for quantum memory. This protocol ensures the integrity of quantum communication through immediate state preparation and encoding, followed by direct transmission to the receiver. This process is detailed as follows[23]:
  1. 1.
[…]
  2. 2.
    Initial Transmission to Alice: Bob sends these qubits to Alice via the quantum channel.
  3. 3.
    Eavesdropping Detection: Alice randomly selects some received qubits for immediate measurement in basis $X$ or $Z$. The results are communicated to Bob through the classical service channel. Bob then verifies if the measured qubits match the initially prepared states. Any discrepancy indicates potential eavesdropping by Eve, causing the process to terminate if the error rate ($e_{e}$) exceeds a predefined threshold. Otherwise, Alice and Bob proceed to estimate the secrecy capacity ($C_{s}$).
  4. 4.
    Message Encoding: Alice encodes the message bits ($M_{k}$) into codewords ($C_{nc}$) using a predetermined coding scheme.
  5. 5.
    Photon Modulation: Alice modulates the remaining qubits by applying either the identity operator ($I$) or the unitary operator ($U$) based on the bit values ’0’ or ’1’ of $C_{nc}$. These modulated photons are then stored in a quantum memory.
  6. 6.
    Return Transmission to Bob: The modulated photons are sent back to Bob through the same quantum channel.
  7. 7.
    Demodulation and Decoding: Bob demodulates the received photons to retrieve the codewords ($C^{\prime}_{nc}$), then decodes these to extract the original message ($Y_{k}$).
Key Features and Advantages
  • Eavesdropping Detection: The immediate measurement of randomly selected qubits allows Alice and Bob to detect any interference by Eve, leveraging the fundamental principles of quantum mechanics where any measurement alters the quantum state.
  • Quantum Memory Utilization: Unlike protocols that operate memory-free, DL04 requires quantum memory, ensuring qubits are stored securely between initial transmission and modulation. This aids in maintaining the integrity and sequence of transmitted qubits.
  • Secrecy Capacity Estimation: The protocol allows for precise estimation of the secrecy capacity ($C_{s}$), which is crucial for determining the security of the communication channel.
  • Two-Way Transmission: The bidirectional flow of qubits between Alice and Bob enhances the robustness of the protocol by providing an additional layer for detecting and mitigating eavesdropping attempts.
@@ II Long-Distance Free-Space QSDC Overview
This section introduces long-distance free-space quantum secure direct communication (QSDC). Our investigation delves into how advanced technologies can surmount existing constraints, rendering QSDC viable for widespread application both in the open air and across space through satellite communication. The process initiates by setting up the quantum state, marking the first phase in encoding quantum information for transmission. Following this, the process involves applying LDPC coding to enhance error correction capabilities, crucial for maintaining the integrity of quantum data over long distances. PAT technologies are integrated to ensure precise alignment and stabilization of the quantum signal. An atmospheric correction step is then applied to mitigate effects like scattering and absorption that can degrade the quantum signal as it traverses the atmosphere. The corrected signal undergoes transmission, sending it across free space, potentially covering vast distances including satellite-to-ground communication paths. Upon reaching the destination, the quantum signal is measured in a critical phase where the encoded quantum information is detected and interpreted. Finally, the process concludes with information decoding, where the quantum data is translated back into classical information for use.
@@ II-A Quantum State Preparation and Encoding
II-A1 Basis Determination ($a_{0}$):
  • If $a_{0}=0$, the basis is computational, with states $\{|0\rangle,|1\rangle\}$.
  • If $a_{0}=1$, the basis is superpositional, where $a_{1}$ further dictates the specific state:
    • $a_{1}=0$: State becomes $\frac{|0\rangle+|1\rangle}{\sqrt{2}}$
    • $a_{1}=1$: State becomes $\frac{|0\rangle-|1\rangle}{\sqrt{2}}$
@@ II-A1 Basis Determination ($a_{0}$):
  • If $a_{2}=0$, the phase states remain as initially encoded.
  • If $a_{2}=1$, the phase states are swapped:
    • $|0\rangle+|1\rangle$ becomes $|0\rangle-|1\rangle$ and vice versa.
@@ II-C Integration of Advanced Quantum Technologies
II-D Security Checking Process
To ensure the security of the QSDC protocol, a detailed security checking process is implemented. After the initial transmission and measurement phases, a verification step is conducted where Alice and Bob compare a subset of their data to check for any discrepancies. This process includes the following steps:
  1. 1.
    Error Rate Estimation: Alice and Bob share a portion of their measurement results to estimate the quantum bit error rate (QBER). If the QBER exceeds a predefined threshold, the communication is deemed insecure, and the process is aborted.
  2. 2.
    Entanglement Verification: By verifying the entanglement of the transmitted quantum states, Alice and Bob can ensure that no eavesdropping has occurred. This involves comparing the measurement results to check for correlations consistent with the expected entangled states.
  3. 3.
    Classical Post-Processing: Any discrepancies identified during the verification steps are corrected through classical post-processing techniques, such as error correction and privacy amplification. This ensures that the final shared data is secure and free from eavesdropping.
By incorporating these security measures, the LF QSDC protocol ensures the secure transmission of quantum information over long distances, making it suitable for practical applications in free-space and satellite-based communication systems.
@@ III-A1 LDPC Code Parameter Optimization
We propose a GNN-enhanced adaptive decoding algorithm for quantum communication, improving LDPC code decoding by adapting to channel changes and enhancing security with decoy states for eavesdropping detection and privacy amplification methods to secure fi- nal keys, significantly increasing transmission reliability and protection.
Standard BP Equation:
@@ III-A3 Security Enhancement Techniques
  • After decoding, apply a hash function $H$ to the key to reduce its length and eliminate partial information[28, 29].
@@ V-A Technical Implementation of LF QSDC
Figure 4: Summarizes the main characteristics of free-space channel and their impacts on optical signal transmission[36, 37]
@@ VIII Acknowledgement
Author Contributions: These authors contributed equally: Yew Kee Wong, Yifan Zhou, Xinlin Zhou, Yan Shing Liang, and Zi Yan Li.
Under the guidance of Yew Kee Wong, Yifan Zhou, Yan Shing Liang, Xinlin Zhou, and Zi Yan Li conducted the research with Yifan Zhou and Yan Shing Liang leading the methodology development. The original draft was primarily written by Yifan Zhou, with significant contributions from Xinlin Zhou, Zi Yan Li, and Yan Shing Liang. Xinlin Zhou also undertook the task of reviewing and editing the manuscript. Yew Kee Wong supervised the project, provided revisions, and managed project administration.
IX Data Availability
The datasets analyzed during the current study are available from the corresponding author on reasonable request.